Consider the circuit. If the probability that each switch is closed is $p$,then find the probability of current flowing through $AB$.

  • A
    $p^2 + p$
  • B
    $p^3 + p - 1$
  • C
    $p^3 + p$
  • D
    $p^2 + p + 1$

Explore More

Similar Questions

The probability that India wins the toss is $3/4$. If it wins the toss,the probability of winning the match is $4/5$,otherwise it is $1/2$. Find the probability that India wins the match.

Difficult
View Solution

$A$ purse contains $4$ copper coins and $3$ silver coins,the second purse contains $6$ copper coins and $2$ silver coins. If a coin is drawn out of one of these purses,then the probability that it is a copper coin is :-

If $A_i$ $(i=1, 2, 3, \ldots, n)$ are $n$ independent events with $P(A_i) = \frac{1}{1+i}$ for each $i$,then the probability that none of $A_i$ occurs is

One bag contains $5$ white and $4$ black balls. Another bag contains $7$ white and $9$ black balls. $A$ ball is transferred from the first bag to the second and then a ball is drawn from the second. The probability that the ball drawn is white,is

Difficult
View Solution

If $P(A_1 \cup A_2) = 1 - P(A_1^c) P(A_2^c)$,where $c$ stands for complement,then the events $A_1$ and $A_2$ are

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo